Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Math

Expert replies
by nayeem bari » Wed Apr 15, 2015 12:40 pm
Rituraj has an aquarium, and he will buy exactly 2 of 5 creatures. The available creatures are: two turtles, a goldfish, a catfish and a sucker fish. It is given that a turtle does not fight with the other, neither does it attack the sucker fish. However the turtle will eat any fish in the aquarium. None of the other creatures fights with one another. What is the probability that if Rituraj randomly chooses two creatures, both of them will live umharmed?

A) 6/25
B) 3/5
C) 5/7
D) 1/2
E) Cannot be determined
Join the discussion
Source: — Problem Solving |

by Brent@GMATPrepNow » Wed Apr 15, 2015 12:52 pm
nayeem bari wrote:Rituraj has an aquarium, and he will buy exactly 2 of 5 creatures. The available creatures are: two turtles, a goldfish, a catfish and a sucker fish. It is given that a turtle does not fight with the other, neither does it attack the sucker fish. However the turtle will eat any fish in the aquarium. None of the other creatures fights with one another. What is the probability that if Rituraj randomly chooses two creatures, both of them will live umharmed?

A) 6/25
B) 3/5
C) 5/7
D) 1/2
E) Cannot be determined
The turtle does not .... attack the sucker fish. However the turtle will eat any fish in the aquarium.
Is that correct?
The turtle won't attack the sucker fish, but WILL eat it?

Rituraj's friend: Well it looks like the sucker fish was eaten last night.
Rituraj: Yes, but thankfully it wasn't attacked.
:-)

Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion

by [email protected] » Wed Apr 15, 2015 7:32 pm
Hi nayeem bari,

Brent brings up an interesting point (will a turtle harm the suckerfish or not). I'm going to assume that a turtle will be able to live with a suckerfish and NOT harm it.

The math behind this question can be done in a couple of ways. I'm going to use the Combination Formula and 'map' out the "harmful" options.

Since there are 5 creatures and we're going to choose 2 of them, the order does NOT matter (which means that we can use the Combination Formula to figure out the total number of options):

5c2 = 5!/(2!3!) = 10 different possible pairs of creatures

Now we have to figure out which pairs will lead to a 'harmful' result:

Turtle 1 and goldfish
Turtle 1 and catfish
Turtle 2 and goldfish
Turtle 2 and catfish

So 4 of the 10 options will lead to a 'harmful' result; this means that 6 out of 10 will have an 'UNharmful' result.

6/10 = 3/5

Final Answer: B

GMAT assassins aren't born, they're made,
Rich
Contact Rich at [email protected]
Image
Join the discussion

by Sagar Joshi » Thu Apr 16, 2015 4:51 am
Hi nayeem bari,

Assuming the turtle does not attack the suckerfish, we have 3 possible outcomes

Outcome#1
Selecting 2 Turtles
i.e 2c2/5c2 = 1/10.

Outcome#2
Selecting 1 Turtle and 1 Suckerfish
2c1 * 1c1 / 5c2 = 1/10 (Remember : 2c1 is for selecting a turtle and 1c1 for suckerfish)

Outcome#3
Selecting 2 creatures other than turtles
3c2/5c2 = 3/10

So final probability = 1/10 + 1/5 + 3/10 = 3/5

Hope it helps.
Join the discussion

by Mehdi2 » Thu Apr 16, 2015 4:22 pm
Hi Guys

Why do you guys treat turtles as two different possible selection? Isn't it like dice problems?

Here is my approach:

total number of outcomes: 2c2 (for turtles)+3c1(turtle and other)+3c2(others) = 7

total safe outcomes: 2c2 (for turtles)+1c1 (super + turtle) + 3c2(others) =5

so the answer will be: 5/7
Join the discussion

by Brent@GMATPrepNow » Fri Apr 17, 2015 6:38 am
Rituraj has an aquarium, and he will buy exactly 2 of 5 creatures. The available creatures are: two turtles, a goldfish, a catfish and a sucker fish. It is given that a turtle does not fight with the other, neither does it attack the sucker fish. However the turtle will eat catfish and goldfish. None of the other creatures fights with one another. What is the probability that if Rituraj randomly chooses two creatures, both of them will live umharmed?

A) 6/25
B) 3/5
C) 5/7
D) 1/2
E) Cannot be determined
As Rich demonstrates, we don't really need to get into a lot of probability/counting concepts here.
Once we recognize that the TOTAL number of possible outcomes = 5C2 = 10, we know that there are, AT MOST, 10 scenarios in which no fish die (or do die, as in Rich's solution). This gives us the freedom to simply list those scenarios.

Here are all of the scenarios in which no fish die:
- Turtle 1 and turtle 2
- Turtle 1 and sucker fish
- Turtle 2 and sucker fish
- Goldfish and sucker fish
- Catfish and sucker fish
- Goldfish and catfish
There are 6 such outcomes

So, P(no fish die) = 6/10 = [spoiler]3/5 = B[/spoiler]

Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion

by Matt@VeritasPrep » Sun Apr 19, 2015 11:21 pm
Another (more probabilistic) way of solving this:

Rituraj's creatures will eat each other if he picks (i) a turtle and (ii) a catfish or a goldfish.

There are two ways in which this could happen. (This next step is very tricky, but is important in probability. When a scenario could happen in more than one order, consider both orders!)

1: Turtle first, unlucky fish second.
2: Unlucky fish first, turtle second.

P(1) = (2/5) * (2/4) = 2/10
P(2) = (2/5) * (2/4) = 2/10

Adding this up, we get 4/10, or 2/5. So there's a 2/5 chance the turtle eats one of the fish, meaning there is a (1 - 2/5), or 3/5 chance that harmony reigns in the aquarium. (Let's hope!)
Join the discussion